Define A@B to be A/B + (A x B). What is the value of 20@(4@2)?
Found 2 solutions by jim_thompson5910, Mathtut:
Answer by jim_thompson5910(28595)

First, find 4@2

A@B = A/B + (A x B) ... Start with the given definition

4@2 = 4/2 + (4 x 2) ... Plug in A=4 and B=2

4@2 = 4/2 + (8) ... Multiply

4@2 = 2 + (8) ... Divide

4@2 = 10 ... Add


So 20@(4@2) simplifies to 20@10



Now find 20@10

A@B = A/B + (A x B) ... Start with the given definition

20@10 = 20/10 + (20 x 10) ... Plug in A=4 and B=2

20@10 = 20/10 + (200) ... Multiply

20@10 = 2 + (200) ... Divide

20@10 = 202 ... Add


So 20@(4@2)=202

Answer by Mathtut(3670)

20@(4@2)
A=20
B=4@2
:
(20/(4@2))+20(4@2)
:
4@2
A=4
B=2
:
(20/((4/2)+(4))(2)+20((4/2)+4(2))
:
(20/10)+20(2+8)
:
2+200=202